Night Time 20140322
- II test of new collimation tables. JH wants to check the preset acquire accuracy with and without the dichroic in.
Telescope pointed to BS9143, mr=14, 62 elev.
Preset acquire (agw on axis) in parallactic mode, position angle 90deg (PE90). The rotator angle is at 84.8deg: not good, JH prefers having PE0.
- New preset acquire with PE0 on the same target, rotator angle is at 3.39deg.
Preset active and guiding on.
Dichroic slided in and new preset acquire authorizing LUCI@ARGOS focal station: we find the star within 3": good.
- dichroic slided out and preset acquire authorizing LUCI@RFBG: we find the star within 1": the test of the new collimation table made by JH is succesful!
- Handover to Laser operation, spotters outside and shutter test performed.
- Start of laser propagation, laser pointing adjustment with MAT, cirrus at 10km make laser beams blobby. On PatCams
we see M2 spider. TN: 20140322_070531
Loaded reconstructor for PatCam
vs LM1 offload: 20140320_063219
, 3 manual iterations.
spot are closer as possible to LGSW FoV
We try to acquire the IM between the PatCam
and the LAS periscopes, other adjustment with LM1.
to periscope IM with 5arcsec movement, reconstructor loaded.
We try to offset the LGS position error to periscopes.
fit is not valid!!!
offset evaluated by hand and periscopes moved by hand: the spot is properly centered on PatCam
but we see no light on PnCCD
(just the Blue pupil).
3 LGS recentered
LAN focus stage set to -100k steps
LM1 moved to recenter just BLUE pupil on PnCCD
Time of flight test: 90us better than 70us
Test of ToF
sweep between 9 and 16km: too fast the change of ToF
. RED and YELLOW pupils change behaviour with sweep. BLUE one stay fixed. We try also a hand sweep. 90km seems the better position.
We sweep also the LAN focus. We put it to -60k.
We close both PI and pupil loops
just pupil loop closed
We try to record the IM between PI and periscope: 20140322_082121
The measure offset are too big. We try a IMX0.1
- periscope gain at 0.9, bad. gain at 0.6. The fit procedure is not working when the spot is in the LGSW FoV
. A quad cell algorithm could be implemented.
- Invalid fit fixed by LB putting 0 to the measured position in case of non converging fit...
- We try to close the pupil and jitter loops
- Sweeo ToF
: 90us choosed by eye, this corresponds to 13.5 km
- We try to apply a static shape to M2: 2um iof Z5
New slope_null with M2 flat: 20140322_090256
Then we apply 10um of Z5 to M2
- we try to seep the LAN focus taking snapshots to measure the FHWM:
LAN @ -180k steps, not good
LAN @ -160k
LAN @ -140k
LAN @ -120k
LAN @ -100k
LAN @ -80k
LAN @ -60k, we recenter the LGS, not good
LAN @ -60k
LAN @ -40k
We recenter the LGS
LAN @ -20k
Now we put LGS on PatCams
and we sweep again:
LAN @ -180k steps, spot on PatCams
LAN @ -160k steps, spot on PatCams
LAN @ -140k steps, spot on PatCams
LAN @ -120k steps, spot on PatCams
LAN @ -100k steps, spot on PatCams
LAN @ -80k steps, spot on PatCams
LAN @ -60k steps, spot on PatCams
Yellow LGS appears a little more out of focus with respect the other 2 LGS...
The roundish spot is at -120k.
Recenter LGS with LM1 and periscope loops.
mix of internal source + LGS, we're vignetting the red LGS
- Switched on FLAO, telescope is collimated on WT10_292 (mr=6.5, 14.5deg from Zenith), AO loop closed and new flat saved for M2 (ARGOS and ARGOS1)
New slope_null with flattened M2: 20140322_120511
snapshot with FLAO on and Jitter Open Loop
- Telescope back to Zenith having set ARGOS1 flat, M2 is back to banana shape...
good snapshot to define subapertures.